### v2

# You Only Sample Once: Taming One-Step Text-To-Image Synthesis by Self-Cooperative Diffusion GANs

Abstract
We introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis.This is achieved by integrating the diffusion process with GANs. Specifically, we smooth the distribution by the denoising generator itself, performing self-cooperative learning. We show that our method can serve as a one-step generation model training from scratch with competitive performance. Moreover, we show that our method can be extended to finetune pre-trained text-to-image diffusion for high-quality one-step text-to-image synthesis even with LoRA fine-tuning. In particular, we provide the first diffusion transformer that can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without explicit training.
1 Introduction
Diffusion models (DMs) [46, 12, 50] have recently emerged as a powerful class of generative models, demonstrating state-of-the-art results in many generative modeling tasks, such as text-to-image [41, 58, 6, 7], text-to-video [4, 13], image editing [9, 5, 33] and controlled generation [61, 34]. However, the generation process of DMs requires iterative denoising, leading to slow generation speed. This exacerbates the computational demands of large-scale DMs. The sluggish generation time, coupled with the intensive computational requirements of large-scale DMs, constitutes a substantial barrier to their practical application and wider adoption.
Sampling from DMs can be regarded as solving a probability flow ordinary differential equation (PF-ODE) [50]. Some previous works [47, 28, 29, 2] focus on developing advanced ODE-solver, for reducing the sampling steps. However, they still require typically 20+ steps to achieve high-quality generation. Another line is to distill models from pre-trained PF-ODEs [49, 27, 30, 31, 42], aiming to predict multi-step solution of PF-ODE solver by one step. Existing methods [30, 31] can achieve reasonable sample quality with 4 to 8 steps. However, it is still challenging to generate high-quality samples with one step.
In contrast, Generative Adversarial Networks (GANs) [8, 39] are naturally built on one-step generation with fast sampling speed. However, it is hard to extend GANs on large-scale datasets due to the serious training challenges, with worse sample quality compared to DMs [43, 17]. In this work, we present a novel approach to combine diffusion process and GANs. Previous works [52, 57, 58] have developed some variants of diffusion-GAN hybrid models. However, they rely on adding noise to smooth the adversarial divergence, yielding indirect learning in the one-step generation. In order to directly learn the one-step generation by adversarial divergence and avoid unstable learning, we propose defining the one-step generation distribution on different levels of corrupted samples by diffusion process. And we regard the one-step generation based on less corrupted samples as ground truth to perform adversarial divergence. This approach can be regarded as a self-cooperative process, as the generator learns from itself. We find the innovative design enables stable training and effective learning for one-step generation. Hence, we named our model YOSO which means You Only Sample Once. Besides, we extend our introduced algorithm to one-step text-to-image generation based on pre-trained models by several principled designs. Thanks to our effective designs, we can fine-tune existing text-to-image pre-trained models (i.e., Stable Diffusion [41] and PixArt- [6]) for high-quality one-step generation (See Figure 1). It is worth noting that we even can develop one-step text-to-image models by Low Rank Adaptation (LoRA) [16] fine-tuning.
Our work presents several significant contributions:
We introduce YOSO, a novel generative model that can generate high-quality images with one-step inference.
We propose several principled and effective training techniques enabling the fine-tuning of pre-trained text-to-image DMs for one-step text-to-image, requiring only ~10 A800 days.
We develop the first one-step diffusion transformer [37] with PixArt-, and the first LoRA enabling one-step generation, both of which will be released soon.
We conduct extensive experiments to demonstrate the effectiveness of the proposed YOSO, including image generation from scratch, text-to-image generation fine-tuning, compatibility with existing image customization, and image controllable modules.
2 Related Works
Text-to-image Diffusion Models. Since Diffusion models have shown a stable training process and are well-suited for scaling up generative models, numerous works have been proposed to extend DMs for text-to-image generation [40, 1, 41, 59]. Latent DMs [41, 38] are widely adopted for high-resolution image and video generation due to their computational efficiency.
Few-step Text-to-image Generation. Recently, several works have been developed for few-step text-to-image generation. The LCM [30] builds upon consistency distillation [49], adapting it to Stable Diffusion (SD) and enabling image generation in 4 steps with acceptable quality. Nevertheless, reducing the sampling steps led to poor generation quality. InstaFlow [27] builds upon the Rectified Flows [26, 25], adapting it to Stable Diffusion, facilitating the generation of text-to-image in a mere single step. Despite this, the fidelity of the generated images is still poor. Moreover, due to the model’s training on fixed data-noise pairings, it loses the flexibility to support variable data-noise combinations, which is detrimental to tasks like image modification or iterative multi-step sampling that can enhance the quality of generated images. Recently, combining DMs and GANs for one-step text-to-image generation has been explored. UFOGen extends DDGANs [52] and SSIDMs [57] to Stable Diffusion by modifying the computation of reconstruction loss from corrupted samples to clean samples. However, it still performs adversarial matching using corrupted samples. ADD [44] introduces a one-step text-to-image generation based on stable diffusion. It follows earlier research by employing a pre-trained image encoder, DINOv2 [35], as the backbone of the discriminator to accelerate the training. However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands.Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training. This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training. In contrast, we perform the adversarial training by replacing real data with self-generated data to smooth the ground truth. Moreover, compared with UFOGen and ADD, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them. Additionally, we extend our method not only to stable diffusion but also to PixArt- [6] which is based on Diffusion Transformer [37]. This demonstrates the wide application of our proposed YOSO.
3 Background
Diffusion Models. Diffusion models [46, 12] define a forward process that gradually transforms samples from data distribution to Gaussian distribution by adding noise in steps with variance schedule : . The corrupted samples can be directly obtained by , where and . The parameterized reversed diffusion process is defined to gradually denoise: . The model can be trained by minimizing the negative ELBO [12, 20]: where is Gaussian posterior derived in [12]. A key assumption in diffusion is that the denoising step size from to is sufficiently small. This assumption ensures the true follows Gaussian distribution, enabling the effectiveness of modeling with Gaussian distribution.
Diffusion-GAN Hybrids. An issue in diffusion models is that the true does not follow Gaussian distribution when the denoising step size is not sufficiently small. Therefore, in order to enable large denoising step size, Diffusion GANs [52] propose to minimize the adversarial divergence between model and :
where and is imposed by a GAN generator. The capability of a GAN-based formulation, enables much larger denoising step sizes (i.e., 4 steps) compared to Gaussian.
4 Method: Self-Cooperative Diffusion GANs
4.1 Revisiting of Diffusion GANs
Diffusion-GAN hybrid models are tailored for training with a large denoising step size. However, a key issue is that they match the generator distribution at corrupted data distribution. This formulation only indirectly learns the and , which are the distribution used for one-step generation, making the learning process less effective.
4.2 Our Design
To enable more effective learning for one-step generation, we propose to directly construct the learning objectives over clean data. We first construct a sequence distribution of clean data as follows:
where is the data distribution, and is the denoising generator. Note that is our denoising generator that predicts clean samples. If the network is parameterized to predict noise, we have .
Given the constructed distribution, we can formulate the optimization objective as follows:
The optimization objective is constructed by the combination of adversarial divergence and KL divergence. Specifically, the adversarial divergence focuses on matching over the distribution level, ensuring the generation quality, and the KL divergence focuses on matching over point level, ensuring the mode cover.
However, it is hard to directly learn the adversarial divergence over clean data distribution, akin to the challenges with GAN training. To tackle these challenges, previous Diffusion GANs [52, 57] have pivoted towards learning the adversarial divergence over the corrupted data distribution. Unfortunately, as analyzed before, such an approach fails to directly match , curtailing the efficacy of one-step generation. Moreover, it also compels the discriminator to fit different levels of noise, leading to limited capability.
Recall that is defined as . The quality of the distribution has two key factors: 1) the ability of the trainable generator ; 2) the information given by .
Hence given the generator fixed, if we increase the information in , it is supposed that we can get a better distribution. In other words, it is highly likely that is superior to . Motivated by the cooperative approach [53, 54, 55, 11] which uses an MCMC revised version of model distribution to learn the generator, we suggest to use as ground truth to learn , which constructs the following training objective:
where denotes stop-gradient operator. This training objective can be regarded as a self-cooperative approach, since the ‘revised’ samples from and samples from are generated by the same network . Note that we only replace data distribution as in the adversarial divergence for smoothing the learning objective, as recent work [32] has found that it is beneficial to learn the generator with a mix of real data and revised data.
In the distribution matching objective above, we apply the non-saturating GAN objective to minimize the adversarial divergence of marginal distributions. And the KL divergence for point matching can be optimized by loss. Hence a tractable training objective is formulated as follows:
where is the discriminator network. We find that the self-cooperative approach is connected with Consistency Models [49]. However, Consistency Models consider the as an approximated ODE solution of to perform a point-to-point match. In contrast, our proposed objective matches and in marginal distribution level, which avoids the approximated error of ODE related to denoising step size.
To further ensure the mode cover of the proposed model, we can add consistency loss to our objective as regularization, which constructs our final training objective:
where is a time-dependent hyper-parameter.
5 Towards One-Step Text-to-Image Synthesis
Since training a text-to-image model from scratch is quite expensive, we suggest using pre-trained text-to-image diffusion models as initialization with Self-Coopeartive Diffusion GANs. In this section, we introduce several principled designs for developing a generation model that enables one-step text-to-image synthesis based on pre-trained DMs.
5.1 Using pre-trained models for Training
Latent Perceptual Loss. Prior research [15, 14, 49] has confirmed the effectiveness of perceptual loss in various domains. Notably, recent studies [27, 49] has found that the LPIPS loss [62] is crucial for obtaining few-step diffusion models with high sample quality. However, a notable drawback is that the LPIPS loss is computed in the data space, which is expensive. In contrast, the popular Latent diffusion models operate in the latent space to reduce computational demands. Hence, using LPIPS loss in training Latent diffusion models is pretty expensive, as it requires not only the computation of LPIPS loss in the data space but also an additional decoding operation. Realizing that the pre-trained Stable Diffusion can serve as an effective feature extractor [56], we suggest using pre-trained Stable Diffusion to perform latent perceptual loss. However, Stable Diffusion is a UNet, whose final layer predicts epsilon with dimensions identical to the data. Hence we propose using the bottleneck layer of the UNet for computing:
where is the latent encoded images by VAE and is the text feature. We note that the benefit of computing latent perceptual loss by stable diffusion is not only the computational efficiency but also the incorporation of text features, which are crucial in text-to-image tasks.
Latent Discriminator. Training GANs for the text-to-image task on large-scale datasets faces serious challenges. Specifically, unlike unconditional generation, the discriminator for the text-to-image task should justify based on both image quality and text alignment. This challenge is more obvious during the initial stage of training. To address this issue, previous pure GANs [18] for text-to-image propose complex learning objectives and require expensive costs for training. The learning of GANs has been shown to benefit from using a pre-trained network as the discriminator. As discussed above, the pre-trained text-to-image diffusion has learned representative features. Hence, we suggest to apply the pre-trained stable diffusion for constructing the latent discriminator. Similar to latent perceptual loss, we only use half UNet for the discriminator followed by a simple predict head. The advantages of the proposed strategy are twofold: 1) we use the informative pre-trained network as initialization; 2) the discriminator is defined over latent space, which is computationally efficient. Unlike previous work that defines a discriminator over data space which required decode latent and backward from decoder, yielding expensive computational cost. By applying the Latent Discriminator, we observe a more stable training process with fast convergence. Note that, expensive R1 regularization is not needed in our approach.
5.2 Fixing the Noise Scheduler
A common issue in diffusion models is that the final corrupted samples are not pure noise. For example, the noise schedule used by the Stable Diffusion makes terminal Signal-to-noise ratio (SNR) far from reaching zero. The corrupted samples at the final timestep follows:
where the terminal SNR is , effectively creates a gap between training and inference. Previous work [24] has only observed that such an issue makes diffusion models unable to produce pure black or white images, which does not hurt sample quality seriously. However, we find that this issue yields serious problems in one-step generation. As shown in Figure 2, there are notable artifacts in one-step generation if we directly sample noise from standard Gaussian distribution. The reason may be that in multi-step generation, the gap can be gradually fixed in sampling, while one-step generation reflects the gap more directly. To address this issue, we provide two simple yet effective solutions.
Informative Prior Initialization (IPI). The non-zero terminal SNR issue is similar to prior hole issues in VAEs [22, 3, 21]. Therefore, we can use informative prior instead of non-informative prior to effectively address this issue. For simplicity, we adopt a learnable Gaussian Distribution , whose optimal formulation is given below:
where and can be efficiently estimated by finite samples, and follows standard Gaussian distribution. The proposed Informative prior initialization effectively addresses the "prior hole issue" in one-step generation. As shown in Figure 2, the artifacts in one-step generation are immediately removed. We note that the performance is achieved with minimal adjustment, enabling the possibility of developing one-step text-to-image generation by LoRa fine-tuning.
Quick Adaption to V-prediction and Zero Terminal SNR. The Informative prior initialization is not a perfect solution, and it suffers from numerical instability when terminal SNR is pretty low. As shown in Figure 3, we fine-tune PixArt- [6] whose terminal SNR is 4e-5 with introduced technique and -prediction, failing in one-step generation. As discussed in [24], the solution should be the combination of v-prediction [42] and zero terminal SNR. However, we found that directly switching to v-prediction and zero terminal SNR is computationally expensive, requiring lots of extra iterations. This is unacceptable in solving large-scale text-to-image with limited computational resource. To address this, we propose an adapt-stage to fine-tune pre-trained diffusion models to v-prediction and zero terminal SNR.
Adapt-stage-I switch to v-prediction : where , , denotes the parameters of the desired v-prediction model, denotes the parameters of the frozen per-trained model with -prediction.
Adapt-stage-II switch to zero terminal SNR : Note that in adapt-stage-II, we change the noise scheduler to zero terminal SNR. But for the teacher models, we keep the noise scheduler as the same. This approach avoids numerical instability issues of -prediction with zero terminal SNR. We note that the zero terminal SNR has lower SNR than the original schedule at each timestep, so the loss formulates an effective objective to construct a zero terminal SNR student by distilling from the teacher.
We notice that this adapt-stage converges rapidly, typically requiring only 1k iterations for training. This enables a quick adaption to v-prediction and zero terminal SNR from pre-trained -prediction models, thereby comprehensively mitigating the non-zero terminal SNR issue in the noise schedule.
6 Experiments
In this section, we evaluate the proposed YOSO for the image synthesis problem. In Section 6.1, we evaluate YOSO on unconditional image generation by training from scratch, and we conduct ablation studies to highlight the effectiveness of our algorithm. Section 6.3 examines YOSO in the context of text-to-image generation by fine-tuning pre-trained PixArt- [6] and Stable Diffusion [41]. In Section 6.4, we show that the YOSO can be used for several downstream applications.
| Model | FID | NFE | Time (s) |
| YOSO | 3.82 | 1 | 0.05 |
| Ablation: | |||
| YOSO w/o consistency loss | 4.41 | 1 | 0.05 |
| YOSO w/o LPIPS loss | 4.63 | 1 | 0.05 |
| YOSO w/ | 10.85 | 1 | 0.05 |
| UFOGen [58] | 45.15 | 1 | 0.05 |
| DDGANs [52] | 3.75 | 4 | 0.21 |
| CT [49] | 8.70 | 1 | - |
| iCT [48] | 2.83 | 1 | - |
| DDPM [12] | 3.21 | 1000 | 80.5 |
| EDM [19] | 2.04 | 36 | - |
| Score SDE (VP) [50] | 2.41 | 2000 | 421.5 |
| Probability Flow (VP) [50] | 3.08 | 140 | 50.9 |
| DDIM [47] | 4.67 | 50 | 4.01 |
| FastDDPM [23] | 3.41 | 50 | 4.01 |
| StyleGAN2 w/o ADA | 8.32 | 1 | 0.04 |
| StyleGAN2 w/ADA | 2.92 | 1 | 0.04 |
6.1 Unconditional Image Generation
Experiment Setting. For the model architectures, we use the same UNet and discriminator architecture as used in DDGANs [52]. We set the denoising steps as four. For the evaluation metric, we choose the Fréchet inception distance (FID) [10]. We train the proposed model on the CIFAR-10 dataset [60]. Following Consistency Models [49], we employ LPIPS loss instead of MSE to compute image distance.
Results. The quantitative results are shown in Table 1. We observe that our method provides competitive performance compared with the best DMs and GANs. Although some variants of DMs achieve better results, they generally require much longer sample time compared to ours. We also find that among variants of GANs, our performance is surpassed only marginally by StyleGAN2 employing adaptive data augmentation. Note that we do not employ data augmentation in experiments. We also observe that the consistency models built upon the diffusion process provide competitive and even better results compared to us. However, we note that the training of consistency models requires tons of computational resources and careful design of the time schedule. As shown in Figure 4, the consistency models require much more computational resources than ours, including more parameters, larger batch size, and more training iterations. To ensure a fair comparison, we re-produce the consistency models by using the same batch size and training iterations as ours. Based on the fair setting, we find that the performance of consistency models significantly decrease, showing way worse performance compared to YOSO. We highlight that the low demand for computational resources enables us to scale the proposed YOSO to large-scale text-to-image datasets.
6.2 Ablation Studies
We provide additional insight into the effectiveness of the proposed YOSO by performing ablation studies to some potential variants on CIFAR-10 dataset.
Effect of Consistency Loss and LPIPS Loss. We evaluate the effect of consistency loss and LPIPS loss in YOSO. We find that the both consistency loss and LPIPS loss can improve the image quality, as indicated by FID. But we notice that if removing consistency loss or LPIPS loss, the generation performance would not decrease significantly. This indicates the effectiveness of our proposed adversarial formulation.
Adversarial Divergence . The key design of our proposed YOSO is smoothing the distribution by using self-generated data instead of real data as ground truth to perform adversarial divergence. We evaluate the variant of using as adversarial divergence. As shown in Table 1, the performance of the variant is significantly worse than YOSO. Moreover, besides sample quality, we find that the training of our proposed YOSO is much more stable than the variant. As shown in Figure 4, the discriminator loss of the variant is not stable and is much smaller than ours. This indicates that our approach can effectively smooth the distribution to make the distributions of fake and real samples closer to each other, making discrimination harder and hence resulting in higher discriminator loss. Furthermore, we observe mode collapse in the training of the variant. We note that a similar case also happened in SDXL-Turbo, a strong one-step text-to-image model, which also performs adversarial divergence directly against real data (See Figure 5). These results indicate the superiority of the proposed approach.
Comparison to UFOGen. Similar to us, UFOGen does not directly perform adversarial divergence against real data either. However, they perform adversarial divergence over corrupted data (i.e., ). We argue that the formulation is less effective to construct one-step generation model since their adversarial divergence only indirectly matches the one-step generation at clean data. We implement the UFOGen on CIFAR-10 with our own. Note that we also perform consistency loss in the UFOGen variant to ensure a fair comparison, as the consistency loss has been found useful in YOSO. As shown in Table 1, UFOGen only obtains 45.15 FID on CIFAR-10, significantly worse than YOSO with 3.82 FID. This further demonstrates the superiority of our proposed approach. Through comparison with different adversarial divergence formulations, we find that our formulated adversarial divergence not only smooths the distribution to stabilize the training but also provides effective learning on one-step generation.
6.3 Text-to-image generation
Experiment Setting. We initialize the GAN generator by pre-trained PixArt- which is a diffusion transformer with 0.6B parameters. We note that 0.6B parameters can be demonstrated as lightweight among text-to-image generation models. For the GAN discriminator, we construct the latent discriminator by pre-trained Stable Diffusion 1.5 using the proposed construction introduced in Section 5.1. We switch the pre-trained PixArt- to v-prediction by the proposed technique introduced in Section 5.1, followed by training on the JourneyDB dataset [36] with resizing as 512 resolution. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We observe the training convergence fast, requiring only around 10 A800 days to be trained.
Evaluation. For image quality, we employ Aesthetic Score (AeS) [45] to evaluate. We also adopt the Human Preference Score v2.1 (HPS v2.1) [51] to evaluate the image-text alignment and human preference. The AeS is computed by an aesthetic score predictor trained on LAION [45] datasets, without considering the image-text alignment. HPS is trained to predict human preference given image-text pairs, considering the image-text alignment and human aesthetic. We mainly compare our model against the open-source state-of-the-art models, i.e., SD-Turbo [44] and SDXL-Turbo [44].
Quantitative Results. The quantitative results are presented in Table 3. To provide a direct comparison with SD-Turbo, we also fine-tune a model based on pre-trained SD 1.5 on JourneyDB with the introduced technique. We would like to highlight that it is better to conduct training on LAION dataset [45] which is the training set used by SD 1.5. However, it is not available when we conducted this research. Fine-tuning SD 1.5 on other datasets suffers from data shift, facing a risk of catastrophic forgetting. As shown in Table 3, both YOSO-SD and YOSO-PixArt- clearly beat the SD-Turbo which is based on SD 2.1, a better SD compared to SD 1.5. It is also observed that YOSO-PixArt- is able to compete with SDXL-Turbo which is a much bigger model with 2.6B parameters.
Qualitative Comparison. To provide more comprehensive insight in understanding the performance of our YOSO, we further provide the qualitative comparison in Figure 5. The comparison shows that: YOSO clearly beats SD-Turbo in terms of both image quality and prompt alignment. Specifically, the samples by SD-Turbo are blurry to some extent. Although SDXL-Turbo with a larger model size improves the sample quality compared to SD-Turbo, we find that the samples by YOSO is sharper than SDXL-Turbo to some extent, and we do better in prompt alignment. For example, in the first row of Figure 5 related to a Pikachu, both SD-Turbo and SDXL-Turbo cannot follow the prompt to generate angry expressions and red eyes, while YOSO follows the prompt more precisely. Moreover, it is important to highlight that SDXL-Turbo seems to have a serious issue in mode collapse. Specifically, the first row and third row of Figure 5 show that samples by SDXL-Turbo are really close to each other. This indicates that SDXL-Turbo performs poorly in mode cover. Overall, based on quantitative and qualitative results, we conclude that our YOSO is better in sample quality, prompt alignment, and mode cover compared with the state-of-the-art one-step text-to-image models (i.e., SD-Turbo and SDXL-Turbo).
6.3.1 Zero-Shot One-Step 1024 Resolution Generation
Due to the computational resource limitation, we trained YOSO on 512 resolution. However, there is a 1024-resolution version of PixArt-, obtained by continuing training on the 512-resolution version of PixArt-. Motivated by the effectiveness of the LoRA combination [31], we suggest constructing a similar combination as follows:
where and means the model weight. We let . As shown in Figure 1(b), the YOSO constructed in this way can generate high-quality images with 1024 resolution. Note that YOSO even changes the predicted objective to v-prediction. The impressive performance indicates the robust generalization ability of our YOSO.
6.3.2 Fine-tuning Stable Diffusion via LoRA
We provide experiments of full fine-tuning PixArt- and SD 1.5 on JourneyDB above. It is worth mentioning that JourneyDB is not in the training set of SD 1.5, and training on it will cause a distribution shift to some extent, which hinders the wide application in downstream tasks. Hence, we further conduct experiments on fine-tuning SD 1.5 via LoRA. As we perform LoRA fine-tuning, it is expected that it will be worse than full fine-tuning, we propose to use SDXL-Turbo to generate data for LoRA fine-tuning, which avoids the risk of data shift to some extent and is economically efficient. We call the model fine-tuned in this way as YOSO-LoRA. We present the quantitative results in Table 4 and qualitative results in Figure 6. It can be observed that YOSO-LoRA archives even better results than LCM with 4 steps and is even comparable to InstaFLow with full fine-tuning. And from Figure 6, YOSO-LoRA has a clearly better sample quality than LCM with 2 steps and InstaFlow. These results indicate that our algorithm enables performing one-step text-to-image synthesis with reasonable quality by LoRA fine-tuning. However, we still observe a little bit of distribution shift compared to stable diffusion in YOSO-LoRA. This issue can only be addressed by training on the LAION dataset [45] or generating training data by SD 1.5 which is expensive.
6.4 Application
One promising and attractive property of text-to-image DMs is that they can be used in multiple downstream tasks as the base model. In this section, we show the capability of our proposed YOSO in various downstream tasks, keeping the unique advantages of text-to-image DMs. Since the relevant applications are mostly developed for stable diffusion, we mainly show the results obtained by YOSO-LoRA in this section.
Image-to-image. As shown in Figure 6, our YOSO-LoRA is capable of performing high-quality image-to-image editing [33] in one step.
Compatibility with ControlNet and Different Base Models. As shown in Figure 6, our YOSO-LoRA is compatible with ControlNet [61], following the condition well. Our YOSO-LoRA is also compatible with different base models (e.g., dreamshaper and realistic vision) fine-tuned from SD 1.5, preserving their style well.
We find that when applying YOSO-LoRA to new base models, it fails to produce reasonable samples in one step. However, YOSO-SD succeeds in doing so when combined with different base models using Equation 8. Unfortunately, the style of the new base models cannot be preserved, as YOSO-SD also alters the style of the original SD 1.5 due to distribution shift in our training set.
7 Conclusion
In summary, we present YOSO, a new generative model enabling high-quality one-step generation. Our novel designs combine the diffusion process and GANs, enabling not only one-step generation training from scratch but also one-step text-to-image generation fine-tuning from pre-trained models. We will release our model to advance the research of text-to-image synthesis.